OpenStudy (anonymous):

How do I simplify √2-1 / √2 + 1? The answer is 3-2√2. But I got 1.. D:

OpenStudy (anonymous):

I tried: √2 - 1 √2-1 √2 - 1 ------ = ----- x ------ √2 + 1 √2 + 1 √2 - 1

OpenStudy (anonymous):

can u be clear \[\frac{\sqrt{2-1}}{\sqrt{2+1}}\] use conjugates \[\huge{\frac{\sqrt{2-1}}{\sqrt{2+1}}*\frac{\sqrt{2-1}}{\sqrt{2-1}} = \frac{(2-1)}{\sqrt{3}} = \frac{1}{\sqrt{3}}}\]

OpenStudy (anonymous):

no it's just: √2 - 1 / √2 + 1

OpenStudy (anonymous):

\[\huge\frac{\sqrt{2-1}}{\sqrt{2-1}}=1\]

OpenStudy (mimi_x3):

\[\large \frac{\sqrt{2}-1}{\sqrt{2}+1} *\frac{\sqrt{2}-1}{\sqrt{2}-1} = \frac{(\sqrt{2}-1)(\sqrt{2}-1)}{2-1} = \frac{3-2\sqrt{2}}{1} =3-2\sqrt{2} \]

OpenStudy (anonymous):

why is the denominator: 2-1?

OpenStudy (anonymous):

and the denominator: 3 - 2√2?

OpenStudy (escherichiarinku):

(a+b)(a-b)=a^2+b^2

OpenStudy (anonymous):

ohh..

OpenStudy (escherichiarinku):

and the nominator is (a-b)(a-b) = (a-b)^2 expanded into a^2-2ab+b^2

OpenStudy (mimi_x3):

because \((\sqrt{a}+\sqrt{b})(\sqrt{a}-\sqrt{b}) = a - b\)

OpenStudy (anonymous):

ohh, thank you! I have to memorize those right?

OpenStudy (mimi_x3):

hm..if you want to lol. You can use FOIL to work it out anyway, might be harder; depends on you xD

OpenStudy (anonymous):

ohh, i see.

OpenStudy (escherichiarinku):

At school we were all forced to memorize such formulas, lol. Well, if you work on enough problems the simplification formulas will be intuitive :D

OpenStudy (anonymous):

ohh, thanks guys!

OpenStudy (anonymous):

Mimi, how did you get 3 - 2√2/1? D:

OpenStudy (anonymous):

'cause don't I have to multiply: √2*√2*√2*-1?

OpenStudy (mimi_x3):

I used FOIL, to do it. Or you can use \[(\sqrt{a}-\sqrt{b})^{2} = a-2\sqrt{ab} +b\] I don't know what you are doing there, sorry.